OpenStudy (anonymous):

What are the coordinates of the x-intercepts of the graph of y = 2x2 + 6x – 20? (–5, 0), (2, 0) (5, 0), (–2, 0) (–4, 0), (10, 0) (4, 0), (–10, 0)

OpenStudy (hunus):

The x intercepts occur when y = 0. So factor 2x^2 + 6x - 20 = 0 2(x^2 + 3x - 10 = 0) x^2 + 3x - 10 = 0

OpenStudy (anonymous):

So what answer would that be?

OpenStudy (hunus):

Well, try factoring x^2 + 3x - 10 = 0

OpenStudy (anonymous):

Oh okay yeah one sec

OpenStudy (anonymous):

(5,0) (-2,0)?

OpenStudy (anonymous):

Oh my bad negative 5 positive 2

OpenStudy (hunus):

Yup :)

OpenStudy (anonymous):

Thanks :)